Abrianna built a bookcase that is 8 feet tall and 6 feet wide. She put a diagonal brace across the back of the bookcase, as shown below. What is the length of the diagonal brace?
Answer by derangoj(14)


a^2 + b^2 = c^2
8^2 + 6^2 = 100
SQRT(100) = 10
the length is 10ft